# Muse: A Multimodal Conversational Recommendation Dataset with Scenario-Grounded User Profiles

### v1

@@ 4.5 Evaluation on Conversational Recommendation Task: Response
Additionally, we randomly sample 100 pairs of zero-shot and finetuned responses from different models and combine them for anonymous manual evaluation as Appendix D.5. As shown in the Table.6, responses generated from fine-tuned models are more favored by users. Interview feedback indicated that evaluators generally find fine-tuned responses better capture users’ key interests within the given context.
@@ A.1 Data Preprocess
A.2 Details of the Open-dialogues
@@
The workflow of the multimodal-open case can be seen in Figure. 6. Specifically, an MLLM (gpt-4o-mini) is employed to analyze outfit requirements based on the user’s target product/scenario, automatically generating descriptions of potential matching items. These descriptions are served as queries to search our product database. Since the retrieved local product may not perfectly match these generated descriptions - a commom occurence - the model performs an additional compatibility assessment. The secondary evaluation determines whether the retrieved local products can effectively coordinate with the user’s target item while fulfilling the intended scenario requirements. When the secondary evaluation is successful (with an approximate success rate of 47%), the conversation adopts an multimodal-open format, in which users express both their scenario-related requirements and outfit-item-driven needs, allowing the rec-assistant to identify suitable product recommendations.
A.3 Architecture of Rewriter
@@
A.4 Multimodal Large Language Model Settings for Synthesizing Data
@@ B.1 Discussion on Data Quality
B.2 Cost Analysis
@@
B.3 Discussions of N-gram Diversity Patterns
The aforementioned analysis reveals that the Distinct-4 scores are significantly low in both the Pearl and Muse (without the Rewriter). This phenomenon can be attributed to the behavior of GPT-series models operating at low temperatures. While these models demonstrate strong instruction-following capabilities, they tend to generate dialogues with highly repetitive sentence patterns. To provide concrete evidence of this pattern repetition, we conduct a detailed analysis of the most frequent 4-grams in each dataset, which offers a clear visualization of this linguistic homogeneity issue.
Figure. 9, 10, and 11 present a comprehensive visualization of the 4-gram distribution in each dataset. The left panels display word clouds of the top 200 4-gram phrases, while the right panels show frequency distributions of the top 10 most recurring phrases. The word cloud visualizations clearly demonstrate that the majority of these 4-grams consist of repetitive sentence structures and fixed phrasal patterns. It substantiates our hypothesis that the low Distinct-4 scores are a direct result of this limited linguistic variability, as the prevalence of standardized sentence constructions inherently reduces 4-gram diversity. The analysis reveals a notable disparity in phrase frequency distributions. In both Pearl and Muse (without the Rewriter), the top 10 4-gram phrases exhibit disproportionately high frequencies relative to the total utterance count. However, following the application of Rewriter, we observe a substantial reduction in the frequency of these recurring phrases in Muse. This decline in repetitive patterns, coupled with the increased 4-gram diversity shown in Table 1, provides compelling evidence that the Rewriter successfully enhances linguistic variability by generating more diverse and sophisticated sentence structures. Additional diversity metrics are shown in Table. 7.
@@ B.3 Discussions of N-gram Diversity Patterns
B.4 Discussion of the Necessity of Images
@@ D.1 Scenario-Grounded User Profile Generator
The steps of the Scenario-Grounded User Profile Generator are divided into two steps as in Figure.  15. The first step is to screen the matched basic scenario, user basic profile, and target product. The second step is to generate the backstory of the user to demonstrate his/her preferences.
@@ D.2 Actions of User
Users exhibit two primary actions: accepting a recommendation and rejecting a recommendation, with their corresponding prompt templates illustrated in Figure.  16. Additionally, users demonstrate two secondary action patterns: chit-chatting and comparing item (find reason for rejecting the item).
@@ D.5 Setting for Manual Evaluation
For dialogue quality evaluation at the utterance-level, we employ a two-phase approach as shown in Figure. 19. In the first phase, we recruit annotators to generate artificial dialogues based on clear task instructions without any external prompts or interventions. In the second phase, we compile a test set containing the dialogue context, newly collected responses, and original dialogues from Muse. These are randomly shuffled to ensure unbiased assessment. We then conduct blind A/B testing with a separate group of annotators to get final results.

### v2

@@ 4.5 Evaluation on Conversational Recommendation Task: Response
Additionally, we randomly sample 100 pairs of zero-shot and finetuned responses from different models and combine them for anonymous manual evaluation as Appendix D.5. As shown in the Table.6, responses generated from fine-tuned models are more favored by users. Interview feedback indicated that evaluators generally find fine-tuned responses better capture users’ key interests.
@@ A.1 Data Preprocess
Figure 7: Distribution of mentioned products in Muse.
A.2 Product Distribution
To gain a deeper understanding of our dataset’s composition, we visualize all the mentioned product types, as shown in the Figure. 7. On the right side of the figure, we display the proportions of the three main categories classified at the broadest level: Clothing, Shoes, and Decoration. (Note: The original dataset labels the third category as Jewelry, but we rename it Decoration after observing that items like wallets are also grouped into this category.) Meanwhile, the left side of the figure illustrates the frequency of the top-10 product types.
A.3 Details of the Open-dialogues
@@
The workflow of the multimodal-open case can be seen in Figure. 6. Specifically, an MLLM (gpt-4o-mini) is employed to analyze outfit requirements based on the user’s target product/scenario, automatically generating descriptions of potential matching items. These descriptions are served as queries to search our product database. Since the retrieved local product may not perfectly match these generated descriptions - a commom occurence - the model performs an additional compatibility assessment. The secondary evaluation determines whether the retrieved local products can effectively coordinate with the user’s target item while fulfilling the intended scenario requirements. When the secondary evaluation is successful (with an approximate success rate of 47%), the conversation adopts an multimodal-open format, in which users express both their scenario-related requirements and outfit-item-driven needs, allowing the rec-assistant to identify suitable recommendations.
A.4 Architecture of Rewriter
@@
A.5 Multimodal Large Language Model Settings for Synthesizing Data
@@ B.1 Discussion on Data Quality
B.2 Extended Manual Verification
In the main text, we use manual verification to perform utterance-level comparisons, where humans generate responses based on the same context and compare them with the responses in Muse. To provide more comprehensive manual comparisons, we introduce an additional comparison experiment here. In this experiment, we collect more human-generated replies and replace selected responses in the existing Muse conversations. Specifically, we randomly select 50 complete conversations from Muse and, for each conversation, replace 4 responses (two from users and two from the rec-assistant) with human-generated replies to create a conversation-level variant. Another user then judges which conversation—original or modified—is better and smoother overall. The results, shown in the Table. 7, show that even under these replacement conditions, the original conversations in Muse are still preferred by human evaluators.
         Data          Muse          Muse (Replace)
         Win Rate          0.68          0.32
Table 7: Comparison of Win Rates between Original Muse and Response-replaced Muse
B.3 Cost Analysis
@@
B.4 Discussions of N-gram Diversity Patterns
The aforementioned analysis reveals that the Distinct-4 scores are significantly low in both the Pearl and Muse (without the Rewriter). This phenomenon can be attributed to the behavior of GPT-series models operating at low temperatures. While these models demonstrate strong instruction-following capabilities, they tend to generate dialogues with highly repetitive sentence patterns. To provide concrete evidence of this pattern repetition, we conduct a detailed analysis of the most frequent 4-grams in each dataset, which offers a clear visualization of this linguistic homogeneity.
Figure. 10, 11, and 12 present a comprehensive visualization of the 4-gram distribution in each dataset. The left panels display word clouds of the top 200 4-gram phrases, while the right panels show frequency distributions of the top 10 most recurring phrases. The word cloud visualizations clearly demonstrate that the majority of these 4-grams consist of repetitive sentence structures and fixed phrasal patterns. It substantiates our hypothesis that the low Distinct-4 scores are a direct result of this limited linguistic variability, as the prevalence of standardized sentence constructions inherently reduces 4-gram diversity. The analysis reveals a notable disparity in phrase frequency distributions. In both Pearl and Muse (without the Rewriter), the top 10 4-gram phrases exhibit disproportionately high frequencies relative to the total utterance count. However, following the application of Rewriter, we observe a substantial reduction in the frequency of these recurring phrases in Muse. This decline in repetitive patterns, coupled with the increased 4-gram diversity shown in Table 1, provides compelling evidence that the Rewriter successfully enhances linguistic variability by generating more diverse and sophisticated sentence structures. Additional diversity metrics are shown in Table. 8.
@@ B.4 Discussions of N-gram Diversity Patterns
B.5 Discussion of the Necessity of Images
@@
B.6 Discussion about the Differences between Muse and Existing Multimodal Datasets
We further elaborate on the differences between Muse and existing multimodal datasets, mainly about SURE and SIMMC Long et al. (2023); Crook et al. (2021); Kottur et al. (2021). These datasets constrain conversations to specific VR scenarios and focus on multimodal interactions centered around recommendation tasks. They primarily emphasize spatial relationships in conversations, such as "Please introduce the red clothes just above the jeans," while providing limited content related to user preferences. Additionally, the products discussed in these conversations are restricted to a small selection available within the VR environment, and these products often lack detailed descriptions. The "conversational recommendation task" described in this paper refers to a conversation driven by user needs and interests. The goal is to identify products that meet user requirements from a large pool of options, based on the interest feedback provided by the user. Each product is accompanied by unique descriptions, including both images and text. This setup aligns with the task definition of traditional recommendation systems, such as collaborative filtering and sequential recommendation tasks, where the objective is to accurately identify products that match user preferences from massive datasets. The task scenario in Muse mirrors real-world situations, such as engaging in a conversation with customer service while shopping online. Therefore, fundamental differences exist between datasets like Muse and SURE-type VR datasets. Muse aligns more closely with the definition of conversational recommendation tasks as outlined in existing datasets Li et al. (2018); Zhou et al. (2020).
@@ D.1 Scenario-Grounded User Profile Generator
The steps of the Scenario-Grounded User Profile Generator are divided into two steps as in Figure.  16. The first step is to screen the matched basic scenario, user basic profile, and target product. The second step is to generate the backstory of the user to demonstrate his/her preferences.
@@ D.2 Actions of User
Users exhibit two primary actions: accepting a recommendation and rejecting a recommendation, with their corresponding prompt templates illustrated in Figure.  17. Additionally, users demonstrate two secondary action patterns: chit-chatting and comparing item (find reason for rejecting the item).
@@ D.5 Setting for Manual Evaluation
For dialogue quality evaluation at the utterance-level, we employ a two-phase approach as shown in Figure. 20 and Figure. 21. In the first phase, we recruit annotators to generate artificial dialogues based on clear task instructions without any external prompts or interventions. In the second phase, we compile a test set containing the dialogue context, newly collected responses, and original dialogues from Muse. These are randomly shuffled to ensure unbiased assessment. We then conduct blind A/B testing with a separate group of annotators to get final results.